# Cornwall-Jackiw-Tomboulis effective field theory to nonuniversal equation of state of an ultracold Bose gas

### v1

@@ B.3 Two-loop
In equation (39), $G\left(k\right)$ is the propagator or Green’s function. By minimizing the CJT effective potential density $V_{\text{eff}}\left[\phi_{0},G\right]$ concerning the elements of the propagator $G\left(k\right)$, we obtain
@@
By optimizing the CJT effective potential density with respect to the order parameter $\phi_{0}$, we can derive the gap equation within the Hartree-Fock (HF) approximation,
@@
$=$ $\frac{1}{8\pi^{2}}\left(\frac{2mM^{2}}{\hbar^{2}}\right)^{\frac{3}{2}}\sqrt{\frac{m}{m^{*}}}\frac{\Gamma\left[1\right]\Gamma\left[\frac{-3}{2}\right]}{\Gamma\left[-\frac{1}{2}\right]}$
@@
Next, we substitute the expression for $M$ from Eq. (56) into Eq. (C) and organize the resulting equation into orders of $\epsilon^{0}$, $\epsilon^{1}$ and $\epsilon^{2}$.

### v2

@@ B.3 Two-loop
In equation (39), $G\left(k\right)$ is the propagator or Green’s function. By minimizing the CJT effective potential $V_{\text{eff}}\left[\phi_{0},G\right]$ concerning the elements of the propagator $G\left(k\right)$, we obtain
@@
By optimizing the CJT effective potential with respect to the order parameter $\phi_{0}$, we can derive the gap equation within the Hartree-Fock (HF) approximation,
@@
$=$ $\frac{1}{8\pi^{2}}\left(\frac{2m^{*}M^{2}}{\hbar^{2}}\right)^{\frac{3}{2}}\sqrt{\frac{m}{m^{*}}}\frac{\Gamma\left[1\right]\Gamma\left[\frac{-3}{2}\right]}{\Gamma\left[-\frac{1}{2}\right]}$
@@
Next, we substitute the expression for $M$ from Eq. (56) into Eq. (55) and organize the resulting equation into orders of $\epsilon^{0}$, $\epsilon^{1}$ and $\epsilon^{2}$.